OpenStudy (anonymous):

Alright lets try a fun one. It costs a family of 8 a total of $209 for general admission tickets to the Mexican City Zoo. How many adult tickets and how many child tickets were purchased? GOGOGO!

OpenStudy (johnweldon1993):

I believe it's Impossible to know without knowing the price of each adult and each child ticket...

OpenStudy (ja1):

Well we don't know 1. How many children/adults and 2. How much the tickets are (either or)

OpenStudy (anonymous):

Adult $33 Child $22

OpenStudy (anonymous):

i was trying to upload a file but internet went dumb on me my bad

OpenStudy (johnweldon1993):

So we know A + C = 8 and 33A + 22C = 209 solve the system...

OpenStudy (johnweldon1993):

So lets solve for adults first C = 8 - A plug that into the second equation 33A + 22(8 - A) = 209 33A + 176 - 22A = 209 33A - 22A = 11A 11A + 176 = 209 -176 -176 11A = 33 --------- 11 11 A = 3 There are 3 adults...how many kids?

OpenStudy (anonymous):

3 adult, 5 children...

OpenStudy (johnweldon1993):

Lets check that 33A + 22C = 209 33(3) + 22C = 209 99 + 22C = 209 -99 -99 22C = 110 ---------- 22 22 C = 5 so yes correct

OpenStudy (anonymous):

nice job im impressed. What property did you just used?

OpenStudy (johnweldon1993):

It is called "substitution"

OpenStudy (anonymous):

by God i could not solve it oh well thanks!

OpenStudy (johnweldon1993):

When you have a system like that....you solve for 1 variable...that way you can substitute that in...and solve for the other variable as I did :) If you need anymore help / explanation feel free to message me :)